# Enhancing Fault Localization Through Ordered Code Analysis with LLM Agents and Self-Reflection

### v1

@@ Abstract.
Locating and fixing software faults is a time-consuming and resource-intensive task in software development. Traditional fault localization methods, such as Spectrum-Based Fault Localization (SBFL), rely on statistical analysis of test coverage data but often suffer from lower accuracy. Learning-based techniques, while more effective, require extensive training data and can be computationally expensive. Recent advancements in Large Language Models (LLMs) offer promising improvements in fault localization by enhancing code comprehension and reasoning. However, these LLM-based techniques still face challenges, including token limitations, degraded performance with long inputs, and difficulties managing large-scale projects with complex systems involving multiple interacting components. To address these issues, we introduce LLM4FL, a novel LLM-agent-based fault localization approach that integrates SBFL rankings with a divide-and-conquer strategy. By dividing large coverage data into manageable groups and employing multiple LLM agents through prompt chaining, LLM4FL navigates the codebase and localizes faults more effectively. The approach also incorporates self-reflection and chain-of-thought reasoning, enabling agents to iteratively generate fixes and re-rank suspicious methods. We evaluated LLM4FL on the Defects4J (V2.0.0) benchmark, comprising 675 real-world faults from 14 open-source Java projects. Our results demonstrate that LLM4FL outperforms AutoFL by 19.27% in Top-1 accuracy and surpasses state-of-the-art supervised techniques such as DeepFL and Grace, all without task-specific training. Additionally, we highlight the impact of coverage splitting and prompt chaining on fault localization performance and show that different method ordering can improve Top-1 accuracy by up to 22%.
@@ 1. Introduction
The process of locating and fixing software faults requires significant time and effort. Research shows that software development teams allocate more than half of their budgets to testing and debugging activities (Hait and Tassey, 2002; Alaboudi and LaToza, 2021). As software systems become increasingly complex, the demand for more efficient and accurate fault localization techniques continues to grow. Hence, to assist developers and reduce debugging costs, researchers have developed various fault localization techniques (Li et al., 2019; Lou et al., 2021; Qian et al., 2023b; Li et al., 2021; Abreu et al., 2009; Sohn and Yoo, 2017). These techniques analyze code coverage and program execution to identify the most likely faulty components, assisting developers in finding the fault.
@@
However, despite the advances in fault localization techniques, many existing approaches still struggle with scalability and precision. Traditional methods, such as Spectrum-Based Fault Localization (SBFL), use statistical analysis to analyze coverage data from passing and failing test cases to rank suspicious code elements (Abreu et al., 2006). While these techniques provide valuable insights, their accuracy is lower. Their reliance on statistical correlations between test failures and code coverage does not always capture the deeper semantic relationships needed for more accurate fault localization (Wong et al., 2016; Xie et al., 2013; Le et al., 2013). To address these issues, recent techniques applied machine learning and deep learning models to improve fault localization (Sohn and Yoo, 2017; Zhang et al., 2019b; Li et al., 2021, 2019). These methods enhance the ranking of suspicious code elements by incorporating additional information like code complexity, text similarity, and historical fault data. Researchers also leveraged models like Graph Neural Networks (GNNs) to represent code structures and achieved state-of-the-art fault localization accuracy (Lou et al., 2021; Qian et al., 2023b; Rafi et al., 2024; Lou et al., 2021). However, these techniques often require extensive training data and significant training time.
@@
Recent advances in Large Language Models (LLMs) have shown great potential for fault localization due to their strong language comprehension and generation capabilities (Abedu et al., 2024; Lin et al., 2024). LLMs trained on extensive programming datasets can understand code structure, interpret error messages, and even suggest fixes for common software bugs (Kang et al., 2024; Wu et al., 2023; Pu et al., 2023; Li et al., 2023). These models, with their ability to analyze and process both natural language and code, present an opportunity to significantly improve traditional fault localization methods by incorporating deeper semantic analysis and context-aware reasoning. Wu et al. (Wu et al., 2023), directly present LLMs with faulty methods or classes with test failure information and ask to locate the issue, which provides valuable insights. Kang et al. (Kang et al., 2024) which operates as an automated fault localization technique that leverages LLMs to localize fault given a single failing test. It focuses on method-level fault localization and provides bug location and also a natural language explanation of why a particular code location is likely to be faulty.
Despite their potential, existing LLM-based fault localization techniques face several challenges. The token limitations of LLMs restrict their ability to effectively process long code files or large sets of code coverage data, which is often required when dealing with large-scale software systems (Hadi et al., 2023; Hou et al., 2023; Wu et al., 2023). Additionally, the performance of LLMs can degrade when applied to complex systems that require the model to reason over multiple interacting components, making it difficult to maintain accuracy and consistency across broader projects (Levy et al., 2024; Liu et al., 2024). Furthermore, current LLM-based techniques have yet to fully explore how these models can be effectively integrated with traditional fault localization techniques to maximize their strengths in a complementary and efficient manner (Wu et al., 2023; Kang et al., 2024; Yang et al., 2024; Bin Murtaza et al., 2024).
In this paper, we propose LLM4FL, an LLM-based fault localization. To address the challenges of analyzing large-scale software projects, where code coverage and complexity often exceed LLM token limits (Wu et al., 2023; Liu et al., 2024), LLM4FL implements a divide-and-conquer strategy. Before applying the divide-and-conquer strategy, we use an SBFL technique to sort the covered methods, building on findings from prior research (Chen et al., 2024) that demonstrate LLM performance improves when the order of instructions is carefully considered. Then, we divide the coverage data into manageable groups that the LLM can process within its token limits. Note that, regardless of the order, LLM4FL eventually analyzes every covered method.
In addition to using a divide-and-conquer strategy, LLM4FL takes inspiration from how human developers debug software. Developers typically analyze multiple types of information, including error messages, stack traces, and code snippets, to incrementally narrow down potential faulty components (Alaboudi and LaToza, 2021; Böhme et al., 2017). LLM4FL emulates this process by utilizing two LLM agents collaborating to iteratively and autonomously navigate the code to locate the faults. LLM4FL implements a Tester and a Debugger Agent, each tasked with specialized tools to assist in the fault localization process. The Tester Agent identifies and prioritizes suspicious methods by analyzing the failing test, and stack traces in different groups of covered methods to list a list of highly suspicious methods. It mimics the developer’s process of investigating suspicious areas in the code by understanding the context of error messages. Meanwhile, the Debugger Agent thoroughly evaluates the given list of candidate methods by navigating the code and ranks them based on its analysis. These two agents communicate through a prompt chaining mechanism that allows them to share insights and build upon each other’s findings.
@@
We evaluated LLM4FL using the Defects4J (V2.0.0) benchmark (Just et al., 2014), which contains 675 real-world faults from 14 open-source Java projects. Our results demonstrate that LLM4FL surpasses LLM-based technique AutoFL (Kang et al., 2024), by achieving 19.27% higher Top-1 accuracy. Additionally, LLM4FL outperforms supervised techniques such as DeepFL (Li et al., 2019) and Grace (Lou et al., 2021), even without task-specific training. We also analyzed the impact of individual components within LLM4FL on fault localization accuracy. Our findings indicate that each component plays a significant role in its performance, with coverage splitting and prompt chaining contributing the most. When these components are removed, accuracy drops considerably, underscoring their importance in managing token limitations and facilitating efficient multi-agent collaboration. Moreover, we examined whether the initial ordering of methods provided to the LLM influences performance. The results reveal that method ordering is important, with a Top-1 accuracy difference of up to 22% when comparing an execution-based ordering and the order provided by DepGraph (Rafi et al., 2024).
@@
  • We introduce LLM4FL, a novel LLM-based fault localization technique that employs a divide-and-conquer strategy. This technique groups large coverage data and ranks the covered methods using an SBFL formula. By utilizing multiple agents and prompt chaining, LLM4FL navigates the code iteratively to effectively identify and localize faults.
  • LLM4FL demonstrates superior performance, surpassing AutoFL (Kang et al., 2024) by 19.27% in Top-1 accuracy. It also outperforms state-of-the-art supervised techniques like DeepFL and Grace, achieving these results without requiring task-specific training.
  • Our analysis of LLM4FL’s components shows that key features like coverage splitting and prompt chaining are essential to its fault localization accuracy. Removing these features leads to significant performance declines, emphasizing their importance in handling token limitations and enabling effective agent collaboration.
  • We further investigate the effect of method ordering on LLM performance. The study reveals that different ordering can enhance fault localization accuracy by up to 22% in Top-1 scores. While LLM4FLDepGraph achieves the highest overall accuracy, LLM4FLOchiai offers a more efficient solution, balancing accuracy gains with lower computational overhead, making it practical for broader use.
In short, we provide a strategy to mitigate the token limitation issues in LLM-based fault localization and highlight the impact of initial method ordering. The findings may help inspire future research to improve LLM-based fault localization for large-scale software projects.
Paper Organization. Section 2 discusses related work. Section 3 describes our technique, LLM4FL. Section 4 presents the experimental results. Section 5 discusses the threats to validity. Section 6 concludes the paper.
@@ 2.1. Background
Large Language Models. Large Language Models (LLMs), primarily built on the transformer architecture (Meta AI, 2024; Brown, 2020; Roziere et al., 2023), have significantly advanced the field of natural language processing (NLP). These LLMs, such as the widely recognized GPT3 model with its 175 billion parameters (Brown, 2020), are trained on diverse text data from various sources, including source code. The training involves self-supervised learning objectives that enable these models to develop a deep understanding of language and generate contextually relevant and semantically coherent text. LLMs have shown substantial capability in tasks that involve complex language comprehension and generation (Abedu et al., 2024; Lin et al., 2024), such as code recognition and generation. Recent research has leveraged LLMs in software engineering tasks, particularly in fault localization (Kang et al., 2024; Qin et al., 2024; Yang et al., 2024), where they assist in identifying the faulty code groups responsible for software errors. One of the key advantages of using LLMs in fault localization is their ability to process both natural language and code without re-training, allowing them to analyze error messages, stack traces, and test case information to infer suspicious methods or code sections in an unsupervised zero-shot setting.
@@
LLM Agents. LLM agents leverage LLMs to autonomously execute tasks described in natural language, making them versatile tools across various domains. LLM agents are artificial intelligence systems that utilize LLMs as their core computational engines to understand questions and generate human-like responses. They leverage functionalities like memory management (Zhou et al., 2023) and tool integration (Xia et al., 2024; Roy et al., 2024) to handle multi-step and complex operations seamlessly. The agents can refine their responses based on feedback, learn from new information, and even interact with other AI agents to collaboratively solve complex tasks (Hong et al., 2024; Qian et al., 2023a; Xu et al., 2023; Lin et al., 2024). Through prompting, agents can be assigned different roles (e.g., a developer or a tester), providing more domain-specific responses that help improve the answer (Hong et al., 2024; White et al., 2024; Shao et al., 2023). As their potential expands, LLM agents play a crucial role in advancing automation and boosting productivity in software development. In this paper, we explore using LLM agents to improve fault localization by emulating developers’ debugging process.
@@ 2.2. Related Work
Learning-based fault localization. Recent efforts have focused on improving SBFL with learning-based methods (Sohn and Yoo, 2017; Zhang et al., 2019b; Li et al., 2021; Li and Zhang, 2017; Li et al., 2019; Zhang et al., 2019a). These approaches use machine learning models like radial basis function networks (Wong et al., 2011), back-propagation networks (Wong and Qi, 2009), and convolutional neural networks (Zhang et al., 2019a; Li et al., 2021; Albawi et al., 2017) to estimate suspiciousness scores based on historical faults. Some techniques, such as FLUCCS (Sohn and Yoo, 2017), combine SBFL scores with metrics like code complexity, while others, like DeepFL (Li et al., 2019) and CombineFL (Zou et al., 2019), merge multiple sources such as spectrum-based and mutation-based data (Moon et al., 2014; Papadakis and Le Traon, 2015; Dutta and Godboley, 2021). Graph neural networks (GNNs) have also been applied to fault localization (Qian et al., 2023b; Lou et al., 2021; Qian et al., 2021; Xu et al., 2020a). Techniques like Grace (Lou et al., 2021) and GNET4FL (Qian et al., 2023b) utilize test coverage and source code structure for improved accuracy, while DepGraph (Rafi et al., 2024) refines these approaches by graph pruning and incorporating code change information, resulting in higher performance with reduced computational demands. Although these learning-based techniques show improved results, they require training data that may not be available to every project, and the training process can be expensive due to model complexity.
LLM-Based Fault Localization. Large Language Models (LLMs), such as GPT-4o (OpenAI, 2024), LLaMA (Meta AI, 2024), and ChatGPT (Achiam et al., 2023), demonstrated remarkable abilities in processing both natural and programming languages. LLMs have shown potential in identifying and fixing errors using program code and error logs (Achiam et al., 2023). However, one of the major challenges that LLMs face in fault localization is the token limitation issue. LLMs are restricted to a fixed number of tokens, ranging from 2,000 to 128,000 (OpenAI, 2024; Meta AI, 2024), which poses difficulties in handling large-scale software projects with long stack traces and extensive codebases. This limitation can lead to incomplete analyses, as critical context might be truncated or lost, forcing models to work with fragmented information, ultimately affecting the overall fault localization performance (Hadi et al., 2023; Hou et al., 2023).
Moreover, LLM-based fault localization techniques often focus on localizing faults within small code snippets due to these context limitations. For example, LLMAO (Yang et al., 2024) uses lightweight bidirectional adapters on LLMs to generate suspiciousness scores for code lines, but only within a limited context of 128 lines of code. Similarly, Wu et al. (Wu et al., 2023) prompt ChatGPT with code and error logs to identify faulty lines, but these methods struggle to scale to larger software projects (Liu et al., 2024). AutoFL (Kang et al., 2024) is an LLM-based fault localization technique that detects faulty code locations and generates natural language explanations for bugs. However, it also faces difficulty handling large-scale projects due to LLM context length constraints and reduced performance on complex bugs requiring deeper repository exploration. To address these challenges, there is a growing need for approaches that enable LLMs to localize faults across entire software projects, ensuring they can handle larger inputs while maintaining accuracy and effectiveness. Thus, we propose LLM4FL, that leverages a divide-and-conquer approach to navigate through codebases and identify faults through multiple-agent collaboration.

### v2

# A Multi-Agent Approach to Fault Localization via Graph-Based Retrieval and Reflexion

@@ Abstract.
Identifying and resolving software faults remains a challenging and resource-intensive process. Traditional fault localization techniques, such as Spectrum-Based Fault Localization (SBFL), leverage statistical analysis of test coverage but often suffer from limited accuracy. While learning-based approaches improve fault localization, they demand extensive training datasets and high computational resources. Recent advances in Large Language Models (LLMs) offer new opportunities by enhancing code understanding and reasoning. However, existing LLM-based fault localization techniques face significant challenges, including token limitations, performance degradation with long inputs, and scalability issues in complex software systems. To overcome these obstacles, we propose LLM4FL, a multi-agent fault localization framework that utilizes three specialized LLM agents. First, the Context Extraction Agent applies an order-sensitive segmentation strategy to partition large coverage data within the LLM’s token limit, analyze failure context, and prioritize failure-related methods. The extracted data is then processed by the Debugger Agent, which employs graph-based retrieval-augmented code navigation to reason about failure causes and rank suspicious methods. Finally, the Reviewer Agent re-evaluates the identified faulty methods using verbal reinforcement learning, engaging in self-criticism and iterative refinement. Evaluated on the Defects4J (V2.0.0) benchmark, which includes 675 faults from 14 Java projects, LLM4FL achieves an 18.55% improvement in Top-1 accuracy over AutoFL and 4.82% over SoapFL. It outperforms supervised techniques such as DeepFL and Grace, all without requiring task-specific training. Furthermore, its coverage segmentation and prompt chaining strategies enhance performance, increasing Top-1 accuracy by up to 22%.
Fault localization, large language model, agent
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
The process of locating and fixing software faults requires significant time and effort. Research shows that software development teams allocate more than half of their budgets to testing and debugging activities (Hait and Tassey, 2002; Alaboudi and LaToza, 2021). As software systems become increasingly complex, the demand for more accurate fault localization techniques grows. Hence, to assist developers and reduce debugging costs, researchers have developed various fault localization techniques (Li et al., 2019; Lou et al., 2021; Qian et al., 2023b; Li et al., 2021; Abreu et al., 2009; Sohn and Yoo, 2017). These techniques analyze code coverage and program execution to identify the most likely faulty code, assisting developers in finding the fault.
@@
Despite the advances in fault localization, many existing techniques still struggle with scalability and precision. Traditional methods, such as Spectrum-Based Fault Localization (SBFL), use statistical analysis to analyze coverage data from passing and failing test cases to rank suspicious code elements (Abreu et al., 2006). While these techniques provide valuable insights, their accuracy is lower. Their reliance on statistical correlations between test failures and code coverage does not always capture the deeper semantic relationships needed for more accurate fault localization (Wong et al., 2016; Xie et al., 2013; Le et al., 2013). Recent techniques applied machine learning (Sohn and Yoo, 2017; Zhang et al., 2019b; Li et al., 2021, 2019) and deep learning models (Lou et al., 2021; Qian et al., 2023b; Rafi et al., 2024; Lou et al., 2021) to address these issues to improve fault localization. These methods enhance the ranking of suspicious code elements by incorporating additional information like code complexity, text similarity, and historical fault data. However, these techniques often require extensive training data that may not be available.
@@
Recent advances in Large Language Models (LLMs) have shown great potential for software fault localization due to their strong language comprehension and generation capabilities (Abedu et al., 2024; Lin et al., 2024b). LLMs trained on extensive programming datasets can understand code structure, interpret error messages, and even suggest bug fixes (Wu et al., 2023; Pu et al., 2023; Li et al., 2023). These models, with their ability to analyze and process both natural language and code, present an opportunity to significantly improve traditional fault localization methods by incorporating deeper semantic analysis and context-aware reasoning.
However, there are several challenges in using LLMs for fault localization. Firstly, LLMs face significant difficulties when analyzing large-scale repository-level data, primarily due to inherent token limitations, as highlighted in recent studies (Hadi et al., 2023; Hou et al., 2023; Wu et al., 2023). Secondly, the performance of LLMs tends to decrease when they are applied to complex systems that require reasoning about multiple interacting components. This complexity can negatively impact the performance of LLMs (Liu et al., 2024; Wu et al., 2023). LLMs may struggle to understand the inter-procedural method call relationships, an essential aspect for understanding how faults propagate across the entire software repository (Chatterjee et al., 2015; Levy et al., 2024; Liu et al., 2024).
In this paper, we propose LLM4FL, a multi-agent LLM-based fault localization technique designed to analyze large-scale software projects at the repository level. LLM4FL leverages a multi-agent framework that emulates developers’ fault localization process. LLM4FL employs three LLM agents, each with a specialized role to enhance fault localization effectiveness. The Context Extraction Agent divides the methods covered by the failing test cases into smaller groups, each within LLM’s token limitation. Then, the agent analyzes the test code and test failure messages to reason about possible failure causes. This helps prioritize the most suspicious covered methods from each group before passing them on to the next agent. Next, the Debugger Agent analyzes each covered method and performs graph-based, retrieval-augmented (Graph-RAG) code navigation. It traverses the prioritized methods’ code iteratively along call dependencies to analyze and rank suspicious methods. Finally, the Reviewer Agent re-ranks the identified fault methods collected from all groups using verbal reinforcement learning (Shinn et al., 2024), criticizing and refining each response to ensure more precise fault localization.
@@
We evaluated LLM4FL using the Defects4J (V2.0.0) benchmark (Just et al., 2014), which contains 675 real-world faults from 14 open-source Java projects. Our results demonstrate that LLM4FL surpasses LLM-based technique AutoFL (Kang et al., 2024) and AgentFL/SoapFL (Qin et al., 2024) by achieving 18.55% and 4.82% higher Top-1 accuracy, respectively. LLM4FL is also the cheapest among the three, costing only \$0.05 per fault. Additionally, LLM4FL outperforms supervised techniques such as DeepFL (Li et al., 2019) and Grace (Lou et al., 2021), even without task-specific training. We also analyzed the impact of individual components within LLM4FL on fault localization accuracy. Our findings indicate that each component plays a significant role in performance. Among these, dividing the covered methods into smaller groups, and Graph-RAG-based code navigation contributes the most. Moreover, we examined whether the initial ordering of methods provided to the LLM influences performance. The results reveal that method ordering is important. Even though LLM eventually visits all methods, the Top-1 accuracy still has a difference of up to 22% when comparing an execution-based ordering and the order provided by DepGraph (Rafi et al., 2024).
@@
  • We introduce LLM4FL, a novel LLM-based fault localization technique that employs a divide-and-conquer strategy. This technique groups large coverage data and ranks the covered methods using an SBFL formula. Using multiple agents and code navigation based on Graph-RAG, LLM4FL analyzes the repository iteratively to identify and localize faults.
  • We conducted an extensive evaluation and LLM4FL demonstrates superior performance, surpassing AutoFL (Kang et al., 2024) by 18.55% and AgentFL/SoapFL (Qin et al., 2024) by 4.82% in Top-1 accuracy. It also outperforms supervised techniques like DeepFL and Grace, achieving these results without requiring task-specific training.
  • Our analysis of LLM4FL’s components shows that key components like dividing methods into smaller groups and code navigation are essential to its fault localization accuracy. Removing these features leads to performance declines, emphasizing their importance in handling token limitations and effective fault analysis.
  • We find that the different ordering on the initial method list passed to LLM can affect fault localization accuracy by up to 22% in Top-1 scores, even though LLM eventually visits all the methods.
In short, we provide a strategy to mitigate the token limitation issues and analyze repository-level data in LLM-based fault localization. We also highlight the impact of initial method ordering for LLM’s input. The findings may help inspire future research on LLM-based fault localization for large-scale software projects.
Paper Organization. Section 2 discusses background and related work. Section 3 describes our technique, LLM4FL. Section 4 presents the experimental results. Section 5 discusses the threats to validity. Section 6 concludes the paper.
Table 1. Comparison of LLM-based fault localization techniques.
Aspect LLM4FL (this work) AutoFL (Kang et al., 2024) SoapFL/AgentFL (Qin et al., 2024) AutoCodeRover (Zhang et al., 2024)
Navigating large code repository The agent autonomously navigates the repository, iteratively following an inter-procedural call graph to retrieve and explore connected methods. Retrieve class and method signatures related to the failure, then let the LLM decide which methods need deeper analysis. Analyze the test failures to identify the most suspicious classes, then iteratively analyze related methods. Start with the file or method names extracted from issue reports, then iteratively search the repository based on called methods in the retrieved source code.
Handling token limitation Divides methods into groups and analyzes each group separately. Provide only class and method signatures to the LLM, then let it decide which methods need deeper analysis. Use LLMs to generate description for classes and methods using their signatures. Then, the LLM decides which classes/methods are relevant to the test failures for further analysis. Iteratively search for relevant classes and methods based on the issue description and prioritize methods ranked by SBFL techniques.
Ranking of faulty methods Hierarchical ranking. First, it finds the suspicious methods and ranks them in each group. Then, it employs verbal reinforcement learning to combine and re-rank the suspicious methods from all groups. Re-run the fault localization process multiple times and do a majority vote on the result. Utilize a multi-round dialogue strategy that allows LLM agents to discuss the code to score and rank faulty methods. Does not rank methods; it selects a method as faulty from the retrieved context and generates a patch.
@@ 2.1. Background
Large Language Models. Large Language Models (LLMs), primarily built on the transformer architecture (Meta AI, 2024; Brown, 2020; Roziere et al., 2023), have significantly advanced the field of natural language processing (NLP). These LLMs, such as the widely recognized GPT3 model with its 175 billion parameters (Brown, 2020), are trained on diverse text data from various sources, including source code. The training involves self-supervised learning objectives that enable these models to develop a deep understanding of language and generate contextually relevant and semantically coherent text. LLMs have shown substantial capability in tasks that involve complex language comprehension and generation (Abedu et al., 2024; Lin et al., 2024b), such as code recognition and generation. Recent research has leveraged LLMs in software engineering tasks, particularly in fault localization (Kang et al., 2024; Qin et al., 2024; Yang et al., 2024), where they assist in identifying the faulty code groups responsible for software errors. One of the key advantages of using LLMs in fault localization is their ability to process both natural language and code without re-training, allowing them to analyze error messages, stack traces, and test case information to infer suspicious methods or code sections in an unsupervised zero-shot setting.
@@
LLM Agents. LLM agents leverage LLMs to autonomously execute tasks described in natural language, making them versatile tools across various domains. LLM agents are artificial intelligence systems that utilize LLMs as their core computational engines to understand questions and generate human-like responses. They leverage functionalities like memory management (Zhou et al., 2023) and tool integration (Xia et al., 2024; Roy et al., 2024) to handle multi-step and complex operations seamlessly. The agents can refine their responses based on feedback, learn from new information, and even interact with other AI agents to collaboratively solve complex tasks (Hong et al., 2024; Qian et al., 2023a; Xu et al., 2023; Lin et al., 2024b). Through prompting, agents can be assigned different roles (e.g., a developer or a tester), providing more domain-specific responses that help improve the answer (Hong et al., 2024; White et al., 2024; Shao et al., 2023).
Recent studies (Shinn et al., 2024; Renze and Guven, 2024; Pan et al., 2025) explore using the agent’s verbal reasoning result to guide iterative improvement (i.e., verbal reinforcement learning), which has shown promising improvement in downstream tasks. In verbal reinforcement learning, LLM agents receive natural language feedback, such as reasoning results or instructions, from other agents as a reward signal. This allows agents to learn and adapt their behavior based on human-like guidance, improving their learning process to solve complex tasks. As the capabilities of large language model (LLM) agents grow, they play an essential role in enhancing automation and increasing productivity in software development. They can assist in code generation (Nijkamp et al., 2022; Lin et al., 2024a; Gu, 2023), debugging (Lee et al., 2024; Kang et al., 2024), test case creation (Huang et al., 2024; Chen et al., 2024b), and automated refactoring (Pomian et al., 2024; Liu et al., 2025), enabling developers to streamline repetitive tasks and focus on higher-level design and problem-solving. Additionally, LLM agents can facilitate collaborative software engineering by acting as intelligent assistants in code reviews, documentation generation, and issue resolution, improving overall development efficiency (Lin et al., 2024a; He et al., 2024). In this paper, we explore using LLM agents to improve fault localization by emulating developers’ debugging process using verbal reinforcement learning.
@@ 2.2. Related Work
Learning-based fault localization. Recent efforts have focused on improving SBFL with learning-based methods (Sohn and Yoo, 2017; Zhang et al., 2019b; Li et al., 2021; Li and Zhang, 2017; Li et al., 2019; Zhang et al., 2019a). These approaches use machine learning models like radial basis function networks (Wong et al., 2011), back-propagation networks (Wong and Qi, 2009), and convolutional neural networks (Zhang et al., 2019a; Li et al., 2021; Albawi et al., 2017) to estimate suspiciousness scores based on historical faults. Some techniques, such as FLUCCS (Sohn and Yoo, 2017), combine SBFL scores with metrics like code complexity, while others, like DeepFL (Li et al., 2019) and CombineFL (Zou et al., 2019), merge multiple sources such as spectrum-based and mutation-based data (Moon et al., 2014; Papadakis and Le Traon, 2015; Dutta and Godboley, 2021). Graph neural networks (GNNs) have also been applied to fault localization (Qian et al., 2023b; Lou et al., 2021; Qian et al., 2021; Xu et al., 2020a). Techniques like Grace (Lou et al., 2021) and GNET4FL (Qian et al., 2023b) utilize test coverage and source code structure for improved accuracy, while DepGraph (Rafi et al., 2024) refines these approaches by graph pruning and incorporating code change information, resulting in higher performance with reduced computational demands. Although these learning-based techniques show improved results, they require training data that may not be available to every project.
LLM-Based Fault Localization. Large Language Models (LLMs), such as GPT-4o (OpenAI, 2024), LLaMA (Meta AI, 2024), and ChatGPT (Achiam et al., 2023), demonstrated remarkable abilities in processing both natural and programming languages. LLMs have shown potential in identifying and fixing errors using program code and error logs (Achiam et al., 2023). Some LLM-based fault localization techniques operate on small code snippets due to LLM’s token limitations. LLMAO (Yang et al., 2024) uses bidirectional adapters to score suspicious lines within a 128-line context. In contrast, Wu et al. (Wu et al., 2023) prompt ChatGPT with code and error logs but struggle to scale to large projects (Liu et al., 2024).
Recent fault localization techniques employ different strategies to navigate large codebases and refine fault localization results (Kang et al., 2024; Zhang et al., 2024; Qin et al., 2024). Each technique consists of key components for retrieving failure-related classes or methods, navigating the code repository, and ranking suspicious methods. However, there are differences in how these techniques design the components. Table 1 provides a summary of the key differences between prior techniques and ours. AutoFL (Kang et al., 2024) retrieves class and method signatures related to the test failure, and prompts the LLM to decide which methods need future analysis (i.e., to examine the code). While this approach ensures the input size is less than the LLM’s token limitation, it does not consider method dependence in the analysis. Instead, it refines rankings by repeatedly running the fault localization process and using a majority voting mechanism to determine the most likely faulty methods. AgentFL/SoapFL (Qin et al., 2024) models fault localization as a structured operating procedure with controlled phases. It uses a document-guided approach, where the system creates (or refines) LLM-generated or developer-provided documentation for classes and methods, then navigates the codebase based on these documents. By iteratively scoring and ranking suspicious methods, it continuously narrows down the fault candidates. It integrates document-guided search and enhanced method documentation to navigate the codebase while using iterative LLM-based scoring and ranking to refine the set of suspicious methods. AutoCodeRover (Zhang et al., 2024) first extracts file/method names from an issue report. Then, based on the names, it searches for corresponding classes and methods. It is an iterative search process and may continue to search for the called methods in the repository based on LLM feedback. AutoCodeRover may also be integrated with SBFL to refine localization and guide patch generation. However, it does not explicitly rank faulty methods; instead, it selects a method from the retrieved context and generates a patch.
In contrast to prior techniques, LLM4FL uses a structured, graph-based approach to analyze the code, enabling the LLM to systematically navigate the repository by considering caller-callee relationships. This eliminates LLM hallucination when searching and retrieving methods in the repository. Additionally, LLM4FL uses a divide-and-conquer strategy to mitigate token limitations by breaking down large methods into smaller, manageable groups, ensuring comprehensive analysis. Finally, LLM4FL refines fault localization results through verbal reinforcement learning, iteratively re-ranking suspicious methods for improved accuracy. We believe these contributions collectively enhance the accuracy and scalability of LLM-based fault localization.